Solve the following quadratic equations by completing the square.
x^2 + 4x - 12 = 0

Answer by algebrahouse.com(1659)

x² + 4x - 12 = 0
(x + 6)(x - 2) = 0 {completing the square would be silly, when you can just factor the left side into two binomials}
x + 6 = 0 or x - 2 = 0 {set each factor equal to 0}
x = -6 or x = 2 {solved each equation for x}

For more help from me, visit: www.algebrahouse.com